I don't know at all and having trouble with these equations I don't know even where to startthey are problems like a X equal 7 x-9
Answer by ankor@dixie-net.com(22740)

X equal 7 x-9
assume the problem is
x = 7x - 9
Add 9 to both sides
x + 9 = 7x
Subtract x from both sides
9 = 7x - x
9 = 6x
divide both sides by 6
9/6 = 1x
1.5 = x
:
:
see if that works in the original equation, replace x with 1.5
1.5 = 7(1.5) - 9
1.5 = 10.5 - 9
1.5 = 1.5; equality reigns, that is a good solution
:
perform math operations on both sides until you get x on one side by itself
that solves the problem. then check your answer in the original equation